Your aunt is planning to invest in a bank CD that will pay 5.50 percent interest semi-annually. If she has $7,000 to invest, how much will she have at the end of four years? (If you solve this problem with algebra round intermediate calculations to 4 decimal places, in all cases round your final answer to the nearest penny.)
Answer by Theo(13342)

f = p * (1+r) ^ n

f is the future value
p is the present value
r is the interest rate per time period
n is the number of time periods.

your time periods are in half years.

4 years is equal to 8 half years.

interest rate of 5.5 percent per year is equal to 2.75% per half year time period.

2.75% is equal to .0275

you divide percent by 100 to get decimal equivalent.

formula of f = p * (1+r) ^ n becomes:

f = 7000 * (1.0275) ^ 8

at the end of 4 years, she will have 8696.6639 rounded to 4 decimal places which you round to 8696.67 if you want it to the nearest penny.